# Which Reasoning Trajectories Teach Students to Reason Better? A Simple Metric of Informative Alignment

### v2

@@ 3.4 Proposed Metric: Rank-Surprisal Ratio
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates greater deviation from dominant patterns and provides more informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
@@ 5.1 Trajectory Selection
Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods. maxmax and minmin indicate maximizing or minimizing the corresponding metric. Model names are abbreviated as Q for Qwen and L for LLaMA. Additional results, including GPQA evaluation, are provided in § C.4.
@@
The trajectory selection task aims to identify the most effective reasoning trajectory from a set of candidates for a given problem or prompt. We adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion. This procedure is repeated for all 5,000 training problems and all student models, yielding student-specific 5,000-trajectory teacher datasets for each selection method. We then fine-tune student models on the constructed datasets and evaluate their post-training reasoning performance on standard math benchmarks, consistent with previous experiments. We compare RSR against a random selection baseline and several previously discussed metrics. For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the corresponding trajectory-level score.
@@ Results
Additional trajectory selection results, including complete scores, further analysis, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
@@ 5.2 Teacher Selection
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation. We consider a realistic low-resource setting where generating full training data for every candidate teacher model is either costly or infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the resulting dataset-level average score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@ A.2 Teacher Trajectory Generation
We release all 33 generated trajectory datasets on Hugging Face. [^3] https://huggingface.co/datasets/Umean/RSR_data
@@ A.3 Teacher Models and Student Models
These teachers produce trajectories that vary in length, informational content, and style. For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed. Concrete trajectory examples from different teachers can be inspected in our Hugging Face datasets referenced above.
For student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B.
@@ A.5 Details of Model Fine-Tuning
For trajectory selection experiments in § 5.1, we fine-tune each selected dataset with three different random seeds and report performance averaged over these runs. For teacher–student distillation experiments in § 2.1, however, each teacher already yields three independent trajectory datasets, so we do not further vary random seeds for fine-tuning.
@@ C.2 Additional Results for Correlation Analysis
Table 22, 23, 24, 25, and 26 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
@@ C.4 Additional Results for Trajectory Selection
C.4.1 Ablation Study on Trajectory Selection
@@
Rank-Surprisal Ratiomin 78.6 28.5 53.2 61.4 34.8 51.3
@@
Table 12: Ablation study of trajectory selection with RSRmin-based variant methods across student models. "Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
Table 12 presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
@@ C.4.2 Additional Evaluation on GPQA and Full Results
The complete trajectory selection results underlying Table 6 are presented in Table 17, 18, 19, 20, and 21.
@@ C.4.3 Analysis of Datasets Selected by RSR
Table 14 shows the data composition of datasets selected by RSR across 11 teacher models. The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students. QwQ-32B is generally preferred across student models, consistent with its stable performance. For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed.
@@ D.3 AI Assistant Usage Statement
Prompt

### v3

@@ 3.4 Proposed Metric: Rank-Surprisal Ratio
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates greater deviation from dominant patterns and provides more informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision. § C.5 presents illustrative examples of RSR measurements.
@@ 5.1 Trajectory Selection
Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods (5k samples each). maxmax and minmin indicate maximizing or minimizing the corresponding metric. Model names are abbreviated as Q for Qwen and L for LLaMA. Additional results, including the GPQA evaluation and the No-Selection baseline, are provided in § C.4.
@@
The trajectory selection task aims to identify the most effective reasoning trajectory from a set of candidates for a given problem or prompt. This aligns with the widespread need for data-efficient training in real-world scenarios, particularly when resources are limited. Specifically, we adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion. This procedure is repeated for all 5,000 training problems and all student models, yielding student-specific 5,000-trajectory teacher datasets for each selection method. We then fine-tune student models on the constructed datasets and evaluate their post-training reasoning performance on standard math benchmarks, consistent with previous experiments. We compare RSR against a random selection baseline and several previously discussed metrics. For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the corresponding trajectory-level score.
@@ Results
Additional trajectory selection results, including complete scores, further analysis, the No-Selection baseline, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
@@ 5.2 Teacher Selection
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation, reflecting a recurring challenge in real-world applications. To better capture practical constraints, we consider a low-resource setting in which generating full training data for every candidate teacher model is either costly or infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, evaluate them using different metrics, and select the teacher model based on the resulting dataset-level average score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@ 7 Conclusion
8 Broader Applicability of RSR
Beyond CoT Data
Although our work focuses solely on reasoning tasks, RSR is not specifically designed for reasoning trajectories (CoTs) and can be applied to general text data. Exploring its behavior and effectiveness beyond reasoning settings remains an interesting direction for future work.
Subset Selection
Beyond trajectory and teacher selection, RSR may also be applicable to subset selection. Subset selection aims to identify a high-quality subset from a dataset composed of different reasoning problems and their corresponding trajectories. Unlike trajectory selection, which compares multiple trajectories for the same problem, subset selection requires cross-problem comparison of RSR values to filter samples across heterogeneous reasoning prompts. This scenario is particularly relevant when only a single trajectory is available per problem, yet improved data efficiency is still desired. Although subset selection introduces additional confounding factors due to variation across problems—making it harder to isolate trajectory–student suitability effects—RSR may still help identify relatively more effective samples across different reasoning problems. Preliminary experiments on internal data provide suggestive evidence for this possibility.
@@ A.2 Teacher Trajectory Generation
We release all 33 generated trajectory datasets along with RSR-selected subsets (§ 5.1) on Hugging Face. [^3] https://huggingface.co/datasets/Umean/RSR_data
@@ A.3 Teacher Models and Student Models
These teachers produce trajectories that vary in length, informational content, and style. For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed. Specific trajectory examples from different teachers are available in our released datasets on Hugging Face and in our case study (§ C.5).
For the student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B. We use base models instead of chat models as students, following the training practice of recent reasoning models. This choice ensures a clean starting point with greater potential [41], allows us to observe more substantial training effects, and avoids stylistic mismatch and overlapping supervision between instruction-tuning data and reasoning trajectories.
@@ A.5 Details of Model Fine-Tuning
For trajectory selection experiments in § 5.1, we fine-tune each selected dataset with three random seeds and report performance averaged over these runs. For distillation experiments in § 2.1, however, each teacher already yields three independent trajectory datasets, so we do not further vary random seeds for fine-tuning.
@@ C.2 Additional Results for Correlation Analysis
Table 23, 24, 25, 26, and 27 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
@@ C.4 Additional Results for Trajectory Selection
C.4.1 RSR-based Trajectory Selection Variants and No-Selection Baseline
@@
Rank-Surprisal Ratiomin (5k) 78.6 28.5 53.2 61.4 34.8 51.3
@@
No Selection (55k combined) 72.7 41.5 54.4 58.9 35.3 52.5
Table 12: Performance of RSR-based trajectory selection variants and a No-Selection baseline trained on a 55k combined trajectory dataset. "Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
Although trajectory selection enables data-efficient training and offers substantial practical value under resource-constrained settings, we are also interested in how it compares with training on all available data without any selection when resources are sufficient. To this end, we construct a “No Selection” baseline. Unlike RSR, which selects 5k trajectories from different teachers, this baseline directly merges 11 trajectory datasets—one per teacher—into a unified training set comprising 55k samples. We do not use the full 165k dataset (i.e., all three trajectory datasets per teacher) for two reasons. First, the computational cost would be prohibitively high. Second, training on multiple highly similar trajectory datasets from the same teacher may reduce data diversity and increase the risk of overfitting. We consider the 55k setting to be sufficiently representative and more consistent with practical use cases.
From Table 12, we observe that the 5k trajectories selected by RSRachieve post-training performance that is comparable to—or even exceeds—the 55k-sample “No Selection” baseline on most student models. An exception is LLaMA-3.1-8B, which may require larger-scale data to develop reasoning ability due to its relatively limited prior exposure to reasoning tasks. In contrast, for stronger students (Qwen-3-14B and Qwen-3-4B), the selected 5k trajectories significantly outperform the unfiltered 55k dataset. This suggests that a small amount of high-quality data is sufficient to effectively activate the reasoning capabilities of student models. Overall, these findings demonstrate that trajectory selection not only improves data efficiency but can also deliver superior performance with substantially fewer resources, further highlighting the practical value of RSR.
Table 12 also presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
@@ C.4.2 Additional Evaluation on GPQA and Full Results
The complete trajectory selection results underlying Table 6 are presented in Table 18, 19, 20, 21, and 22.
@@ C.4.3 Analysis of Datasets Selected by RSR
Table 14 shows the data composition of datasets selected by RSR across 11 teacher models. The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students. QwQ-32B is generally preferred across student models, consistent with its stable performance. For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed. We release the RSR-selected datasets for five student models on Hugging Face.
@@ C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
C.5 What Does RSR Prefer? A Token-Level Case Study
To better understand how Rank-Surprisal Ratio operates in practice and what kinds of trajectories it prefers, we conduct a token-level case study comparing trajectories generated by different teacher models.
[…]
As shown, tokens in trajectories from GPT-OSS-20B often exhibit excessively high rank values, suggesting that these trajectories are relatively unfamiliar to the student model and misaligned with the student’s current behavior. This may stem from the fast-paced reasoning flow and uncommon phrasing in GPT-OSS-20B trajectories, which make them harder for the student to follow. Meanwhile, tokens from Nemotron-Super frequently have very low surprisals, indicating that these trajectories closely resemble the student’s own generations and therefore provide limited informativeness. This may be attributed to the relatively conventional reasoning flow and limited analytical depth of Nemotron-Super’s trajectories. As a result, tokens from both models tend to receive high token-level RSR values, implying less effective supervision for student training. By comparison, trajectories from QwQ-32B exhibit a moderate level of comprehensibility and depth. Their tokens show relatively lower rank values alongside higher surprisals, striking a better balance between alignment and informativeness. Consequently, they yield lower token-level RSR values and are preferred under the RSR criterion. These observations are consistent with the distillation results (Table 1) and the subsequent analysis presented in the main body of the paper.
We also observe that token-level RSR may become numerically unstable when extremely low surprisal yields inflated RSR values (e.g., 100+). This motivates our surprisal-weighted averaging for trajectory-level RSR (§ 3.4), which stabilizes the metric and improves its effectiveness.
Additional trajectory examples from 11 teachers are available in our released datasets on Hugging Face.
[…]
Table 17: A case study with visualization of token-level rank, surprisal, and RSR values (with rank clipping) across trajectories generated by GPT-OSS-20B, Nemotron-Super-49B-v1.5, and QwQ-32B, as measured by the Qwen-2.5-7B student. Each token is color-coded according to its RSR value, where darker red indicates lower values (stronger preference under RSR). GPT-OSS-20B exhibits uncommon phrasing and high rank values; Nemotron-Super exhibits conventional reasoning flow and low surprisal; QwQ-32B strikes a more favorable balance between informativeness and alignment, yielding lower token-level RSR values. Trajectory-level RSR is defined as the surprisal-weighted average of token-level RSR values with rank clipping (§ 3.4).
@@ D.3 AI Assistant Usage Statement
Evaluation prompt for LLM-judged quality assessment